# DDBot: Differentiable Physics-based Digging Robot for Unknown Granular Materials

### v3

@@ I INTRODUCTION
The first challenge in developing a practical granular material manipulation system lies in the accurate and efficient prediction of contact dynamics, which is crucial for optimising robotic manipulation solutions [1]. Advances in simulation techniques have significantly progressed rigid object manipulation in the last few decades. In recent years, numerical modelling techniques, such as the material point method (MPM) and its variants, have greatly improved the fidelity and efficiency of simulations of non-rigid materials [2, 3]. MPM models non-rigid materials as particle collections, calculating dynamics based on physical laws on the particles and a background Eulerian grid, achieving superior accuracy compared to data-driven approaches [4]. However, it remains unintuitive to tune the values of granular material properties, such as Young’s modulus, Poisson’s ratio, density, and friction angle, to match the real-world dynamics, hindering the development of high-precision manipulation solutions. While differentiable physics-based system identification (DPSI) has demonstrated efficiency for deformable objects, like plasticine [5], its feasibility on granular materials remains unexplored.
@@
Secondly, there is a lack of investigation of how state-of-the-art planning and control methods perform in such tasks. The most popular learning-based method for robotic manipulation in recent years is arguably deep reinforcement learning (DRL), which has proved successful in many robotic tasks [1]. DRL methods seek to learn a policy that outputs actions to maximise a certain reward function via (mostly) randomised exploration [6]. However, DRL is notoriously data-hungry, and can be easily stuck in local minima due to limited positive feedback. DRL also struggles to generalise to unseen scenarios without fine-tuning or retraining. On the other hand, model predictive control (MPC) is another set of common methods for robotic planning and control [7]. MPC typically applies stochastic sampling or evolutionary methods to optimise a distribution of the manipulation motion with a specified cost function (or fitness function). Despite its successes in various domains, MPC methods tend to evolve solutions based on a cost function that evaluates the system’s final state, lacking consideration of the physical dynamics. It ignores the connection between the manipulation skills and the resultant manipulation outcomes, leading to poor local minima and poor interpretability of the found solutions. In addition, it remains unclear how it may perform in such granular manipulation tasks.
@@
Recently, an emerging set of robotic manipulation methods that rely on differentiable physics-based simulators has shown promising results in generating manipulation solutions for non-rigid materials [8, 9]. These methods combine the realism and accuracy of the laws of physics (e.g., MPM), and the automated differentiation mechanism and parallel computing technique of the simulator (e.g., Torch, TaiChi) [10, 11] to calculate physically meaningful gradients of the manipulation trajectory with respect to a loss function. Though researchers have demonstrated the feasibility of direct gradient-based trajectory optimisation with non-rigid materials that are modelled by the fixed corotated elastic energy and the von Mises plastic criterion [12, 3], no work has been done to study the feasibility of gradient-based methods on granular materials, which are more accurately modelled by the St. Venant-Kirchoff (SVK) elastic energy and the Drucker-Prager (DP) plastic yield criterion [13]. In addition, first-order gradient descent methods, despite their efficiency, are sensitive to numerical instability, particularly in large systems with long trajectories and rugged loss landscapes [14]. Thus, there is a need to study the feasibility of gradient-based optimisation for granular material manipulation.
@@
First, to the best of our knowledge, this is the first attempt to develop an efficient and high-fidelity differentiable simulator specifically designed for granular material manipulation. It supports efficient forward dynamics computation to train data-driven methods, such as DRL, and evaluate non-learning methods, such as MPC. It also enables efficient automatic differentiation to support gradient-based optimisation methods, such as RMSProp [15].
@@
Figure 1: The DDBot (Differentiable Digging Robot) system overview. DDBot has five modules. 1) A robot platform (orange) that performs real-world manipulation and collects real-world surface point clouds of the granular materials. The five small images visualise a real-world digging trajectory that matches the trajectory optimised in simulation shown in the green box. 2) A differentiable physics simulator (green) that simulates the digging skill trajectory given the physics and skill parameters. 3) A differentiable loss function (pink) that compares the geometrical differences between the simulated and real manipulation results. 4) A gradient-based optimisation algorithm (yellow) that updates the physics or skill parameters. 5) The Robot Operating System (ROS) communication protocol and the MoveIt! motion planning library [16] that generates a velocity trajectory from the skill parameters to control the real robot.
@@ II-A Robotic soil manipulation tasks
A vast majority of robotic digging research focuses on soil excavation tasks in the construction context, where a large amount of soil, sand, and rocks are to be removed [17, 18, 19, 20]. An early work presented an autonomous truck-loading system that searches for a starting location of the bucket of a given digging motion to fill the bucket with soils [17]. More recently, methods have been proposed to tackle such excavation tasks with uncertain soil properties and unpredictable soil-bucket interaction forces via hierarchical optimisation [18] and reinforcement learning [19]. Another work also studied the sim-to-real gap for soil loading tasks and demonstrated the possibility of reaching low sim-to-real differences with the discrete element method (DEM) for soils with known and uniform physical properties [20].
In agriculture, heavy machines play a key role in large-scale ploughing, seed-sowing, and harvesting tasks but are typically equipped with tailored soil-digging tools and carried out in a large farming field by a human operator [21], while agricultural robots for planting or seed-sowing are mostly designed as large vehicles operating in large outdoor fields [22, 23]. In terms of indoor and small-scale soil-related manipulations, such as greenhouses and household gardening tasks, not much effort has been spent on automated soil manipulation with precision, flexibility, and adaptability.
Beyond agriculture, soil manipulation is also required to be automated for soil sampling tasks, either for space exploration [24] or determining farming field soil properties [25]. Though these operations are done in a much more constrained space, they are typically carried out by specialised tools, such as drills, that do not consider changing soil properties nor require flexible and high-precision control.
@@ II-B Robotic granular material manipulation methods
After several decades, robotic manipulation methods have evolved from hand-crafted motions and simplified model-based motion planning [26] to data-driven (learning-based) approaches [1] and differentiable physics-based trajectory optimisation [7, 11]. Our focus is on granular material manipulation, whose numbers of dimensions cause classic motion planning methods to struggle, and only a few works in recent years have dealt with such materials.
For granular materials, effort has been devoted to simplifying the representation and dynamics. [*]zhang2020learning [27] trained a reinforcement learning (RL) policy to shape granular materials on a table represented by at most $200$ particles in a simulator powered by position-based dynamics (PBD). Due to the costs of high-fidelity simulations, [*]mateo2020manipulation [28] studied a granular material removal task by simplifying the dynamics into a 2D shape control problem represented by a Gaussian mixture model and optimising the control velocity in the task space to achieve a target shape. [*]tuomainen2022manipulation [29] employed a graph neural network to estimate the particle states of granular material being poured out from a cup and plan trajectories for different desired pouring outcomes. Their training data was, however, collected using Taichi and MPM-based simulation with about $1400$ particles. [*]xue2023neural [30] proposed to avoid the costly computation of the physics-based interactions by using a fully convolutional neural network to achieve differentiable rendering and thus gradient-based trajectory optimisation for tabletop granular material gathering tasks. [*]zhu2023fewshot [31] proposed a meta-learning approach to predict the scooped volume of different granular materials of a scoop action from RGB-D image patches and used Bayesian optimisation to find the parameters of a simplified scoop motion that maximises the scooped volume. Another recent work also presented a framework to train RL policies for granular material transportation tasks using demonstrations collected by gradient-based trajectory optimisation in a differentiable fluid simulation environment to bypass gradient explosion [14].
@@ III-A Problem definition
Actions: An action of the robot manipulator at each step here is a 6D displacement vector in the Cartesian space (i.e., end-effector tip frame) $a=[\Delta x,\Delta y,\Delta z,\Delta a,\Delta b,\Delta c]\in\mathcal{A}\subset\mathbb{R}^{6}$, i.e., it controls the amount of translational changes in the $x$, $y$, $z$ directions and the rotational changes about these three axes. In the real world, motion planning (MOVEIT! and RRT connect [32, 33, 16]) are used to compute a velocity trajectory from a series of Cartesian space waypoints (a trajectory of actions) to control the actual robot arm. In the simulation, the Cartesian displacement for each global step is divided evenly by the number of simulation substeps and applied to the pose of the end-effector before collision detection at each substep.
@@
System dynamics: The dynamics function (or transition probability distribution in stochastic cases) determines the next state of the system given the current state and a given action. Obtaining a perfect function to model such real-world dynamics accurately is nearly impossible. However, it can be approximated using numerical modelling methods. In our work, we adopt the moving least squares material point methods (MLS-MPM) to simulate granular materials [4], coupled with the St. Venant-Kirchhoff (SVK) constitutive model for hyperelasticity and the Drucker-Prager (DP) yield criterion for sand/soil plasticity [34, 13].
MPM represents the soil/sand body as a collection of particles with positions, velocities, deformation gradients, and their material properties, including Young’s modulus, Poisson’s ratio, density, and sand friction angle. It efficiently integrates the system’s velocity over time and handles collision and fracture on a background Eulerian grid [13]. MLS-MPM adopts a novel weak form discretisation of the governing equations and replaces the shape functions in the force computation with MLS approximators, achieving faster and more realistic simulation and two-way coupling with rigid objects that classic MPM fails [4].
For the sake of clarity, we present a simplified procedure of MLS-MPM, the key equations of the constitutive models (hyperelasticity and plasticity) and a brief description of the collision contact process below. Please refer to [13, 4, 5] for detailed illustrations of the dynamic model.
As an explicit simulation algorithm, MLS-MPM advances simulation time in a global/local time-stepping process [4]. In other words, for the $t$-th simulation step ($0<t\leq T$) with a simulation stepsize $\Delta t$, there are $N$ substeps with a sub-stepsize $\Delta t_{sup}=\Delta t/N$. Denote $\boldsymbol{x}^{n+1}_{p}=f_{sub}^{t,n}(\boldsymbol{x}^{n}_{p},a_{t},\boldsymbol{\Theta},\boldsymbol{g},\Delta t_{sup})$ as the function of the $n$-th substep of the $t$-th step, it can be summarised as the list of consecutive functions shown in Algorithm 1.
@@ 10. For each particle $p:$
Please refer to [13] for details of the three conditions. With $\hat{S}_{p}$, the deformation gradient after plasticity is $F_{p}^{\prime}=U_{p}\hat{S}_{p}V^{T}_{p}$ and the Cauchy stress $\sigma_{p}=\frac{1}{det(F^{\prime}_{p})}\bm{P}_{p}^{\prime}{F^{\prime}_{p}}^{T}$, where $\bm{P}_{p}^{\prime}$ is computed by substituting $\hat{S}_{p}$ into Eq. 4, and $\det(F^{\prime}_{p})=1$ is imposed to enforce incompressibility.
Given an action (Cartesian space displacement) at time $t$, the 6D Cartesian pose of the manipulation (agent) is updated by evenly distributing the action over the $N$ substeps and adding it to the manipulator pose at each substep (line 7 in Algorithm 1). The collision behaviours between the particles/grid and the manipulator are modelled by a signed distance field-based collision detection mechanism and the Columb frictional contact model, following the practices of previous works [4, 5].
@@ III-D Gradient-based skill optimisation
Many methods have been developed for such constrained optimisation problems, yet gradient-descent methods are the most attractive due to faster convergence under the condition of the whole system being differentiable (smooth) [35]. We are seeking to compute $\partial r/\partial\boldsymbol{\theta}$, and apply first-order gradient-descent algorithms to iteratively update $\boldsymbol{\theta}$ towards local minima. Three key components need to be differentiable (smooth) in computing the gradients of the skill parameters via the chain rule, i.e., the loss (reward) function $r$, the skill-to-action mapping $\bar{a}(\boldsymbol{\theta})$ and the dynamics function $f(o,\bar{a})$:
@@
The differentiability of the dynamics: Here, we discuss the second term of the right-hand side of Eq 34. As introduced in subsection III-A, the dynamics function for soil/sand-rigid-body interactions is simulated by the MLS-MPM method [4]. The gradient chain of the $t$-th simulation step comprising $N$ substeps from the resultant observation $o_{t+1}$ to the action $a_{t}$ can be written as follows:
@@
Algorithm 1 indicates that, as long as the collision function $f_{collision}$, the grid-to-particle function $f_{g2p}$, the agent position updating function $f_{move\_agent}$, the particle-to-grid function $f_{p2g}$ are differentiable over the particle positions and the agent’s actions, one can obtain $\partial o_{t+1}/\partial a_{t}$ with Eq. 35 using the chain rule. The derivatives of these functions have been provided in [36]. Notice that the constitutive model, $f_{constitutive}$, needs not to be differentiable if only the gradients of actions are of interest. However, our work also seeks to perform system identification that optimises the physics parameters for simulating soil/sand with gradient-based optimisation (see subsection IV-B). Therefore, we further provide proof of the differentiability of the constitutive model in Appendix B.

### v4

@@ I INTRODUCTION
The first challenge in developing a practical granular material manipulation system lies in the accurate and efficient prediction of contact dynamics, which is crucial for optimising robotic manipulation solutions [undef]. Advances in simulation techniques have significantly progressed rigid object manipulation in the last few decades. In recent years, numerical modelling techniques, such as the material point method (MPM) and its variants, have greatly improved the fidelity and efficiency of simulations of non-rigid materials [undefa, undefb]. MPM models non-rigid materials as particle collections, calculating dynamics based on physical laws on the particles and a background Eulerian grid, achieving superior accuracy compared to data-driven approaches [undefc]. However, it remains unintuitive to tune the values of granular material properties, such as Young’s modulus, Poisson’s ratio, density, and friction angle, to match the real-world dynamics, hindering the development of high-precision manipulation solutions. While differentiable physics-based system identification (DPSI) has demonstrated efficiency for deformable objects, like plasticine [undefd], its feasibility on granular materials remains unexplored.
@@
Secondly, there is a lack of investigation of how state-of-the-art planning and control methods perform in such tasks. The most popular learning-based method for robotic manipulation in recent years is arguably deep reinforcement learning (DRL), which has proved successful in many robotic tasks [undef]. DRL methods seek to learn a policy that outputs actions to maximise a certain reward function via (mostly) randomised exploration [undefe]. However, DRL is notoriously data-hungry, and can be easily stuck in local minima due to limited positive feedback. DRL also struggles to generalise to unseen scenarios without fine-tuning or retraining. On the other hand, model predictive control (MPC) is another set of common methods for robotic planning and control [undeff]. MPC typically applies stochastic sampling or evolutionary methods to optimise a distribution of the manipulation motion with a specified cost function (or fitness function). Despite its successes in various domains, MPC methods tend to evolve solutions based on a cost function that evaluates the system’s final state, lacking consideration of the physical dynamics. It ignores the connection between the manipulation skills and the resultant manipulation outcomes, leading to poor local minima and poor interpretability of the found solutions. In addition, it remains unclear how it may perform in such granular manipulation tasks.
@@
Recently, an emerging set of robotic manipulation methods that rely on differentiable physics-based simulators has shown promising results in generating manipulation solutions for non-rigid materials [undefg, undefh]. These methods combine the realism and accuracy of the laws of physics (e.g., MPM), and the automated differentiation mechanism and parallel computing technique of the simulator (e.g., Torch, TaiChi) [undefi, undefj] to calculate physically meaningful gradients of the manipulation trajectory with respect to a loss function. Though researchers have demonstrated the feasibility of direct gradient-based trajectory optimisation with non-rigid materials that are modelled by the fixed corotated elastic energy and the von Mises plastic criterion [undefk, undefb], no work has been done to study the feasibility of gradient-based methods on granular materials, which are more accurately modelled by the St. Venant-Kirchoff (SVK) elastic energy and the Drucker-Prager (DP) plastic yield criterion [undefl]. In addition, first-order gradient descent methods, despite their efficiency, are sensitive to numerical instability, particularly in large systems with long trajectories and rugged loss landscapes [undefm]. Thus, there is a need to study the feasibility of gradient-based optimisation for granular material manipulation.
@@
First, to the best of our knowledge, this is the first attempt to develop an efficient and high-fidelity differentiable simulator specifically designed for granular material manipulation. It supports efficient forward dynamics computation to train data-driven methods, such as DRL, and evaluate non-learning methods, such as MPC. It also enables efficient automatic differentiation to support gradient-based optimisation methods, such as RMSProp [undefn].
@@
Figure 1: The DDBot (Differentiable Digging Robot) system overview. DDBot has five modules. 1) A robot platform (orange) that performs real-world manipulation and collects real-world surface point clouds of the granular materials. The five small images visualise a real-world digging trajectory that matches the trajectory optimised in simulation shown in the green box. 2) A differentiable physics simulator (green) that simulates the digging skill trajectory given the physics and skill parameters. 3) A differentiable loss function (pink) that compares the geometrical differences between the simulated and real manipulation results. 4) A gradient-based optimisation algorithm (yellow) that updates the physics or skill parameters. 5) The Robot Operating System (ROS) communication protocol and the MoveIt! motion planning library [undefo] that generates a velocity trajectory from the skill parameters to control the real robot.
@@ II-A Robotic soil manipulation tasks
A vast majority of robotic digging research focuses on soil excavation tasks in the construction context, where a large amount of soil, sand, and rocks are to be removed [undefp, undefq, undefr, undefs]. An early work presented an autonomous truck-loading system that searches for a starting location of the bucket of a given digging motion to fill the bucket with soils [undefp]. More recently, methods have been proposed to tackle such excavation tasks with uncertain soil properties and unpredictable soil-bucket interaction forces via hierarchical optimisation [undefq] and reinforcement learning [undefr]. Another work also studied the sim-to-real gap for soil loading tasks and demonstrated the possibility of reaching low sim-to-real differences with the discrete element method (DEM) for soils with known and uniform physical properties [undefs].
In agriculture, heavy machines play a key role in large-scale ploughing, seed-sowing, and harvesting tasks but are typically equipped with tailored soil-digging tools and carried out in a large farming field by a human operator [undeft], while agricultural robots for planting or seed-sowing are mostly designed as large vehicles operating in large outdoor fields [undefu, undefv]. In terms of indoor and small-scale soil-related manipulations, such as greenhouses and household gardening tasks, not much effort has been spent on automated soil manipulation with precision, flexibility, and adaptability.
Beyond agriculture, soil manipulation is also required to be automated for soil sampling tasks, either for space exploration [undefw] or determining farming field soil properties [undefx]. Though these operations are done in a much more constrained space, they are typically carried out by specialised tools, such as drills, that do not consider changing soil properties nor require flexible and high-precision control.
@@ II-B Robotic granular material manipulation methods
After several decades, robotic manipulation methods have evolved from hand-crafted motions and simplified model-based motion planning [undefy] to data-driven (learning-based) approaches [undef] and differentiable physics-based trajectory optimisation [undeff, undefj]. Our focus is on granular material manipulation, whose numbers of dimensions cause classic motion planning methods to struggle, and only a few works in recent years have dealt with such materials.
For granular materials, effort has been devoted to simplifying the representation and dynamics. [*]zhang2020learning [undefz] trained a reinforcement learning (RL) policy to shape granular materials on a table represented by at most $200$ particles in a simulator powered by position-based dynamics (PBD). Due to the costs of high-fidelity simulations, [*]mateo2020manipulation [undefaa] studied a granular material removal task by simplifying the dynamics into a 2D shape control problem represented by a Gaussian mixture model and optimising the control velocity in the task space to achieve a target shape. [*]tuomainen2022manipulation [undefab] employed a graph neural network to estimate the particle states of granular material being poured out from a cup and plan trajectories for different desired pouring outcomes. Their training data was, however, collected using Taichi and MPM-based simulation with about $1400$ particles. [*]xue2023neural [undefac] proposed to avoid the costly computation of the physics-based interactions by using a fully convolutional neural network to achieve differentiable rendering and thus gradient-based trajectory optimisation for tabletop granular material gathering tasks. [*]zhu2023fewshot [undefad] proposed a meta-learning approach to predict the scooped volume of different granular materials of a scoop action from RGB-D image patches and used Bayesian optimisation to find the parameters of a simplified scoop motion that maximises the scooped volume. Another recent work also presented a framework to train RL policies for granular material transportation tasks using demonstrations collected by gradient-based trajectory optimisation in a differentiable fluid simulation environment to bypass gradient explosion [undefm].
@@ III-A Problem definition
Actions: An action of the robot manipulator at each step here is a 6D displacement vector in the Cartesian space (i.e., end-effector tip frame) $a=[\Delta x,\Delta y,\Delta z,\Delta a,\Delta b,\Delta c]\in\mathcal{A}\subset\mathbb{R}^{6}$, i.e., it controls the amount of translational changes in the $x$, $y$, $z$ directions and the rotational changes about these three axes. In the real world, motion planning (MOVEIT! and RRT connect [undefae, undefaf, undefo]) are used to compute a velocity trajectory from a series of Cartesian space waypoints (a trajectory of actions) to control the actual robot arm. In the simulation, the Cartesian displacement for each global step is divided evenly by the number of simulation substeps and applied to the pose of the end-effector before collision detection at each substep.
@@
System dynamics: The dynamics function (or transition probability distribution in stochastic cases) determines the next state of the system given the current state and a given action. Obtaining a perfect function to model such real-world dynamics accurately is nearly impossible. However, it can be approximated using numerical modelling methods. In our work, we adopt the moving least squares material point methods (MLS-MPM) to simulate granular materials [undefc], coupled with the St. Venant-Kirchhoff (SVK) constitutive model for hyperelasticity and the Drucker-Prager (DP) yield criterion for sand/soil plasticity [undefag, undefl].
MPM represents the soil/sand body as a collection of particles with positions, velocities, deformation gradients, and their material properties, including Young’s modulus, Poisson’s ratio, density, and sand friction angle. It efficiently integrates the system’s velocity over time and handles collision and fracture on a background Eulerian grid [undefl]. MLS-MPM adopts a novel weak form discretisation of the governing equations and replaces the shape functions in the force computation with MLS approximators, achieving faster and more realistic simulation and two-way coupling with rigid objects that classic MPM fails [undefc].
For the sake of clarity, we present a simplified procedure of MLS-MPM, the key equations of the constitutive models (hyperelasticity and plasticity) and a brief description of the collision contact process below. Please refer to [undefl, undefc, undefd] for detailed illustrations of the dynamic model.
As an explicit simulation algorithm, MLS-MPM advances simulation time in a global/local time-stepping process [undefc]. In other words, for the $t$-th simulation step ($0<t\leq T$) with a simulation stepsize $\Delta t$, there are $N$ substeps with a sub-stepsize $\Delta t_{sup}=\Delta t/N$. Denote $\boldsymbol{x}^{n+1}_{p}=f_{sub}^{t,n}(\boldsymbol{x}^{n}_{p},a_{t},\boldsymbol{\Theta},\boldsymbol{g},\Delta t_{sup})$ as the function of the $n$-th substep of the $t$-th step, it can be summarised as the list of consecutive functions shown in Algorithm 1.
@@ 10. For each particle $p:$
Please refer to [undefl] for details of the three conditions. With $\hat{S}_{p}$, the deformation gradient after plasticity is $F_{p}^{\prime}=U_{p}\hat{S}_{p}V^{T}_{p}$ and the Cauchy stress $\sigma_{p}=\frac{1}{det(F^{\prime}_{p})}\bm{P}_{p}^{\prime}{F^{\prime}_{p}}^{T}$, where $\bm{P}_{p}^{\prime}$ is computed by substituting $\hat{S}_{p}$ into Eq. 4, and $\det(F^{\prime}_{p})=1$ is imposed to enforce incompressibility.
Given an action (Cartesian space displacement) at time $t$, the 6D Cartesian pose of the manipulation (agent) is updated by evenly distributing the action over the $N$ substeps and adding it to the manipulator pose at each substep (line 7 in Algorithm 1). The collision behaviours between the particles/grid and the manipulator are modelled by a signed distance field-based collision detection mechanism and the Columb frictional contact model, following the practices of previous works [undefc, undefd].
@@ III-D Gradient-based skill optimisation
Many methods have been developed for such constrained optimisation problems, yet gradient-descent methods are the most attractive due to faster convergence under the condition of the whole system being differentiable (smooth) [undefah]. We are seeking to compute $\partial r/\partial\boldsymbol{\theta}$, and apply first-order gradient-descent algorithms to iteratively update $\boldsymbol{\theta}$ towards local minima. Three key components need to be differentiable (smooth) in computing the gradients of the skill parameters via the chain rule, i.e., the loss (reward) function $r$, the skill-to-action mapping $\bar{a}(\boldsymbol{\theta})$ and the dynamics function $f(o,\bar{a})$:
@@
The differentiability of the dynamics: Here, we discuss the second term of the right-hand side of Eq 34. As introduced in subsection III-A, the dynamics function for soil/sand-rigid-body interactions is simulated by the MLS-MPM method [undefc]. The gradient chain of the $t$-th simulation step comprising $N$ substeps from the resultant observation $o_{t+1}$ to the action $a_{t}$ can be written as follows:
@@
Algorithm 1 indicates that, as long as the collision function $f_{collision}$, the grid-to-particle function $f_{g2p}$, the agent position updating function $f_{move\_agent}$, the particle-to-grid function $f_{p2g}$ are differentiable over the particle positions and the agent’s actions, one can obtain $\partial o_{t+1}/\partial a_{t}$ with Eq. 35 using the chain rule. The derivatives of these functions have been provided in [undefai]. Notice that the constitutive model, $f_{constitutive}$, needs not to be differentiable if only the gradients of actions are of interest. However, our work also seeks to perform system identification that optimises the physics parameters for simulating soil/sand with gradient-based optimisation (see subsection IV-B). Therefore, we further provide proof of the differentiability of the constitutive model in Appendix B.